Question 3 20 pts You are managing the investments of a pension fund. You have to make a payment to your customers in 6 monthplease answer
Answer #1

Future Value at the end of 5 months = 2,000,000/(1 + 0.05/12) = $1,991,701

Calculating Monthly Deposit,

Using TVM Calculation,

PMT = [PV = 0, FV = 1,991,701, N = 5, I = 0.05/12]

PMT = $395,034.50